OpenStudy (anonymous):

You have a gas sample that is currently at an unknown volume. You need the gas to expand to fill a balloon that has a capacity of 50 L. You know its pressure is 1.05 atm and that you have 1.90 mol of the gas. What temperature does the gas need to be at to fill the balloon? A:337˚C B:337 K B:64 K

OpenStudy (jfraser):

solve \[P*V = n*R*T\] for Temperature, and plug in the numbers. The value of R is\[0.0821\frac{L*atm}{mol*K}\]

OpenStudy (anonymous):

SO? B?

OpenStudy (anonymous):

HELP!!!!!!!!!!!

OpenStudy (jfraser):

that's what i get